An aqueous solution is prepared by dissolving 100 mg/L ofCaCl2 and 230 mg/L of Na2SO4 in water. Assuming that these salts dissociate completely,and that the dissociated ions do not undergo any additional reactions, calculate(a) the molar concentration of each ionic species;(b) the ionic strength of the solution;(c) the activity coefficient for each species (with the Davies equation);(d) the activity of each species
Answer #1

(a) The molar masses of CaCl2 and Na2SO4 are 111 g/mol and 142 g/mol respectively.

The number of moles of CaCl2 are \frac {100mg \times 1g}{111g/mol \times 1000mg}= 0.000901 mol

Thus [Ca^{2+}]= 0.000901 mol/L

[Cl^-]=2 \times 0.000901 mol/L =0.00180 mol/L

The number of moles of Na2SO4 are \frac {230mg \times 1g} {142 g/mol \times 1000mg}=0.00162mol

Hence, [SO_4^{2-}]=0.00162mol/L

[Na^+]= 2 \times 0.00162mol/L = 0.00324 mol/L

(b) The ionic strength of the solution is

I = 0.000901 \times (2)^2 + 0.00180 \times (1)^2 + 0.00324 \times (1)^2 + 0.00162 \times (2)^2 = 0.0151

(c) The activity coefficent fro Na2SO4 will be same as the activity coefficient fro CaCl2. It will be

-log f = 0.5 z1z2(\frac {\sqrt {I}}{1+\sqrt{I}}-0.15I)

-log f = 0.5 \times 2 (-1) (\frac {\sqrt {0.0151}}{1+\sqrt {0.0151}}-0.15 \times 0.0151)=0.1072

f=1.28

(d) The activity of Ca2+ ion is 1.28 \times 0.000901= 0.00115

The activity of Cl- ion is 1.28 \times 0.00180 = 0.00230

The activity of Na+ ion is 1.28 \times 0.00324= 0.00415

The activity of SO42- ion is 1.28 \times 0.00162 = 0.00207